# The Boundaries of Tractability in Hierarchical Task Network Planning

### v1

@@ 4 Primitive Task Networks
Let $k$ be the number of states and $w$ the generalized partial order width of the input instance. Set $U$ to be the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. The algorithm is similar to the one presented for Theorem 13. First, test whether Reduction Rule R0 can be applied, and if yes we use it to terminate. Otherwise, decompose $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$. This can be done in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin (2003). Let $A={\{a_{1},\ldots,a_{|A|}\}}$. Let $E={\{e_{1},\ldots,e_{|E|}\}}$ denote the set of equivalence classes among all actions. For each state $s$ in the state transition graph, let $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ be a boolean variable that is TRUE if and only if there is an executable task sequence in $tn$ starting from $s_{0}$ and ending in state $s$ that employs exactly the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and contains exactly $r_{i}$ tasks from $U$ with action equivalence class $e_{i}$, for all $i\in[|E|]$.
We compute this variable for all possible assignments in a dynamic programming manner. Initialize all values to FALSE except for $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$, which is initialized to TRUE. Whenever a variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ that was FALSE before is set to TRUE, also perform the following update step. Let $T^{\prime}\subseteq T$ be such that $t\in T^{\prime}$ if, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Consider all the states $s^{\prime}$ that are adjacent to $s$ in ${\cal G}$. If there is $j\in[w]$ such that the $(h_{j}+1)^{\text{th}}$ task of the chain $c_{j}$ is in $T^{\prime}$, has action $a_{i}$, and $a_{i}$ connects $s$ and $s^{\prime}$ in ${\cal G}$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ to TRUE. This corresponds to taking the next task from a chain and adding it to a given sequence. Another update rule corresponds to adding a task from $U$ instead. If there is an action equivalence class $e_{i}$ that connects $s$ and $s^{\prime}$ in ${\cal G}$, then define the following. Let $T_{i}={\{t\in T\mid\alpha(t)\in e_{i}\}}$ and let $v_{i}$ denote the number of tasks in $V\cap T_{i}$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. If $r_{i}-v_{i}<|T_{i}\cap U|$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{i-1},r_{i}+1,r_{i+1},\ldots,r_{|E|}]$ to TRUE. Once no more variables are updated and set to TRUE, the computation is complete.
@@
The algorithm sets all variables correctly by the same reasoning as in the proof of Theorem 13. For any $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ and action $a\in A$, let $c(a)$ denote the number of tasks in $V$ with action $a$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. Then, the algorithm returns that it is a yes-instance if and only if there is $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ that is TRUE such that, for each action equivalence class $e_{i}\in E$, we have $\sum_{a\in e_{i}}\max{\{0,\min(a)-c(a)\}}\leq r_{i}$, and, for each action $a\in A$, there are at least $\min(a)-c(a)$ tasks of action $a$ in $U$.
Suppose there is such a TRUE variable. Then, there is an executable task sequence that uses $c(a)$ tasks in $V$ with action $a$ for each $a\in A$, and uses $r_{i}$ tasks from $U$ with action in equivalence class $e_{i}$ for all $i\in[|E|]$. As all tasks in $U$ are isolated, two tasks from $U$ can be used interchangeably in a task sequence if they are from the same action equivalence class. Tasks in the sequence can be replaced by tasks from $U$ of the same equivalence class such that there are at least $\min(a)-c(a)$ tasks with action $a$ from $U$ in the sequence for all $a\in A$. Then, the sequence is a solution to Action Executability.
For the other direction, suppose there is a task sequence that solves the Action Executability instance. Let it result in some state $s$, let it use the first $h_{j}$ tasks of each chain $c_{j}$, and let it employ $r_{i}$ tasks of equivalence class $e_{i}$ for all $i\in[|E|]$. Then, the variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ is TRUE. Further, as the sequence contains at least $\min(a)$ tasks of each action $a\in A$, and only $c(a)$ of these are tasks in $V$, it contains at least $\min(a)-c(a)$ tasks of action $a$ from $U$. In particular, this also gives $\sum_{a\in e_{i}}(\max{\{0,\min(a)-c(a)\}})\leq r_{i}$, so the algorithm correctly returns that it is a yes-instance.
@@
Computing the equivalence classes takes $\mathcal{O}\!\left(|{\cal I}|^{3}\right)$ time. As there are $|E|\leq(k+1)^{k}$ equivalence classes, at most $\mathcal{O}\!\left(|T|^{w}\cdot k\cdot|T|^{(k+1)^{k}}\right)$ variables are considered. For each of these, computing all $v_{i}$ takes $\mathcal{O}\!\left(|T|\right)$ time by iterating through all chains time and it can update at most $w+|E|$ variables. Further, it takes $\mathcal{O}\!\left(|T|\right)$ time to check whether a variable yields a yes-instance. Hence, the algorithm takes ${\mathcal{O}\!\left(|T|^{w+(k+1)^{k}}\cdot k\cdot(|T||+w+|E|)+|V|^{2.5}+|\prec|\right)}={\mathcal{O}\!\left(|T|^{w+(k+1)^{k}+1}\cdot k\right)}$ time. ∎
5 Hierarchical Task Networks
Considering HTNs with compound tasks adds another, seemingly opaque layer of difficulty to the considered planning problems. In this section, we show that—at least in terms of classical complexity—we can cleanly characterize the jump from primitive to compound through three measures on compound task networks: we obtain an algorithmic meta-theorem that allows to lift polynomial-time solvability from primitive to compound task networks, if all three of these measures are bounded. Moreover, this result is tight in the sense that none of the three measures can be dropped from the meta-theorem.
The three measures we need are ${C_{\#}}$, ${C_{s}}$, and ${C_{d}}$:
Definition 5.1.
For an HTN planning problem with initial task network $tn=(T,\prec^{+},\alpha)$, we refer to the number of compound tasks in $tn$ by ${C_{\#}}=|{\{t\in T\mid\alpha(t)\in{\cal C}\}}|$. The maximum size of a network that a task can be decomposed into is ${C_{s}}=\max{\{|T_{m}|\mid(c,(T_{m},\prec^{+}_{m},\alpha_{m}))\in M\}}$. The maximum depth $C_{d}$ of a network is defined recursively: ${C_{d}}=0$ for primitive task networks, and a compound task network has ${C_{d}}=i$ if $i$ is the smallest integer such that every decomposition of all compound tasks in the initial network $T$ results in a new network where ${C_{d}}\leq i-1$.
Lemma 15.
A task network $tn=(T,\prec^{+},\alpha)$ can be decomposed into at most ${C_{c}}^{\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}}$ pairwise non-isomorphic primitive task networks $tn^{\prime}=(T^{\prime},{{\prec^{+}}}^{\prime},\alpha^{\prime})$. For each of these, $|T^{\prime}|\leq|T|+{C_{\#}}\cdot({C_{s}}^{{C_{d}}}-1)$.
Proof.
pairwise non-isomorphic task networks. ∎
Having defined our measures of interest, we proceed to formalize the set of problems our meta-theorem will capture. Intuitively, the aim here is to formalize the set of all problems which deal with compound HTNs by decomposing them.
Definition 5.2.
Let PR be an arbitrary HTN planning problem. Then, PR is decomposable if and only if, for all non-primitive task networks $tn$, PR on $tn$ is a yes-instance if and only if there is a primitive task network $tn^{\prime}$ such that $tn$ can be decomposed into $tn^{\prime}$ and $tn^{\prime}$ is a yes-instance of PR.
Note that all problems discussed in this work are decomposable. Next, in order to capture the general notions of “tractability” for various HTN problems on networks where certain measures are bounded, we need to slightly restrict the measures in question, to avoid entirely degenerate measures.
Definition 5.3.
Let $\kappa$ be any numerical measure of HTNs, i.e., a mapping that assigns each HTN a non-negative number. We say that $\kappa$ is stable if there exists a computable function $f$ with the following property: for each primitive HTN $tn$ and each primitive HTN $tn^{\prime}$ obtained from $tn$ by adding or removing a single task, $\kappa(tn^{\prime})\leq f(\kappa(tn))$.
In particular, all measures considered here are stable. The number of states in the state transition graph only depends on the initial state $s_{0}$ and the set of actions $A$, so it does not change when applying decomposition methods. Moreover, adding an element into the network can only increase the generalized partial order width (as well as all the other structural measures mentioned in the manuscript) by at most $1$.
Theorem 16.
Let PR be a decomposable HTN planning problem and $\kappa$ a stable measure. Assume PR is polynomial-time solvable when restricted to primitive networks of constant $\kappa$. Then, PR is polynomial-time solvable when restricted to networks of constant $\kappa$, ${C_{d}},{C_{\#}},{C_{s}}$.
Proof.
Let ${\cal I}$ have task network $tn=(T,\prec^{+},\alpha)$. If there are no compound tasks in $T$, i.e., ${C_{d}}=0$, the statement immediately holds by applying an algorithm for primitive networks. Thus, assume ${C_{d}}\geq 1$. Then, PR in $tn$ can be solved by listing all the primitive instances $tn$ can be decomposed into and solving PR in each of these. By Lemma 15, there are at most ${C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\leq|{\cal I}|^{g({C_{d}},{C_{\#}},{C_{s}})}$ such instances ${\cal I}_{p}$ for $g({C_{d}},{C_{\#}},{C_{s}})=\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}$ and, for each of these, $|{\cal I}_{p}|\leq|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ for some computable function $f^{\prime\prime}$. Further, for each task network $tn_{p}$ of an instance ${\cal I}_{p}$, we have $\kappa(tn_{p})\leq f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})$, for some computable function $f^{\prime\prime\prime}$ as $\kappa$ is stable and $tn_{p}$ is obtained from $tn$ by removing ${C_{\#}}$ compound tasks and adding $f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ primitive tasks. Thus, if primitive instances ${\cal I}_{p}$ can be solved in time at most $|{\cal I}_{p}|^{f(\kappa(tn_{p}))}$, then the total runtime is
$\mathcal{O}\Big{(}(|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}}))^{f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))}$
$\cdot{C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\Big{)}$
$={\mathcal{O}\!\left(|{\cal I}|^{f^{\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})}\right)},$
where $f^{\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})=g({C_{d}},{C_{\#}},{C_{s}})+\\ f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})^{f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))}$. ∎
We complement Theorem 16 by providing lower bounds which show that bounding $C_{d}$, $C_{\#}$, and $C_{s}$ is necessary in order to obtain the result. In particular, we show that if any one of these measures is not bounded by a constant, then some of the problems considered in this article become NP-hard on compound networks, even for settings that were easily solvable on primitive networks. We note that there are several other existing reductions showing the hardness that derives from decomposing non-primitive networks, see, e.g., Erol et al. (1996); Behnke et al. (2015). However, our reductions are stronger in the sense that they bound all but one of the parameters $C_{d}$, $C_{\#}$, $C_{s}$.
Recall that, by Theorem 11, Plan Validation is solvable in polynomial time on primitive networks of constant generalized partial order width. Moreover, as per Theorem 10, the same also holds for Action Executability when additionally bounding the number of states. In contrast, in Theorems 17, 18 and 19, we prove Plan Validation (and also Action Executability) to be NP-hard on general networks even if:
  • any two of $C_{d}$, $C_{\#}$, $C_{s}$, are constant,
  • all possible decompositions of the network have a generalized partial order width of 00,
  • there is only a single state, and
  • $C_{c}$ is also constant.
Theorem 17.
Plan Validation and Action Executability are NP-hard even if ${C_{c}}=7$, ${C_{d}}=1$, ${C_{s}}=2$, $\prec^{+}=\emptyset$, and $\operatorname{prec}(a)$=$\operatorname{del}(a)$=$\operatorname{add}(a)$$\emptyset$ for every $a\in A$.
Proof.
We prove hardness by reducing from 3-SAT-(2,2), an NP-hard 3-SAT variant, where among all clauses each variable appears exactly 4 times: 2 times as a positive literal and 2 times as a negated literal Darmann and Döcker (2021).
We construct an HTN planning instance where the set of actions $A={\{x_{1},\overline{x_{1}},\ldots,x_{n},\overline{x_{n}}\}}$ contains an action for each literal. The initial task network contains a compound task $t_{x}$ for each variable $x$ and a compound task $t_{c}$ for each clause. A variable compound task $t_{x}$ can be decomposed into either a task network containing two tasks with action $x$ or a task network containing two tasks with action $\overline{x}$. Decomposing $t_{x}$ into the first (second, respectively) one of these is interpreted as setting $x$ to TRUE (FALSE, respectively). A clause compound task $t_{c}$ where the clause $c$ contains literals $\ell_{1},\ell_{2},\ell_{3}$ can be decomposed into one of 7 subnetworks, one with no tasks at all, one with a single task with action $\ell_{1},\ell_{2},$ or $\ell_{3}$, or one with two tasks with actions $\ell_{1}$ and $\ell_{2}$, actions $\ell_{1}$ and $\ell_{3}$, or actions $\ell_{2}$ and $\ell_{3}$. Decomposing a clause task is interpreted as acknowledging that exactly the described literals are FALSE. This gives an instance with $\prec^{+}=\emptyset$. For Plan Validation, let the input plan $p$ be any sequence that contains each literal exactly twice. For Action Executability, let $S$ contain each literal exactly twice.
If we are given a yes-instance of 3-SAT-(2,2), consider a satisfying variable assignment and decompose the compound tasks accordingly using the interpretations given above. Then, in the resulting primitive task network, each literal appears exactly twice as follows. If a variable $x$ is set to TRUE, then $t_{x}$ creates exactly 2 tasks with action $x$ and there is no clause task creating a task with action $x$. At the same time, the literal $\overline{x}$ appears in exactly 2 clauses and, as it is FALSE, each of these clause tasks creates a task with action $\overline{x}$. The case in which the variable $x$ is set to FALSE is analogous. As the tasks are executable in any order, this yields a yes-instance for Plan Validation and Action Executability.
For the other direction, suppose we have a yes-instance for Plan Validation or Action Executability. Then, the constructed HTN can be decomposed into a primitive task network with an executable plan such that each literal appears at least twice. Consider the variable assignment given by interpreting the decomposition of the variable tasks. Towards a contradiction, assume that there is a clause $c$ with literals $\ell_{1},\ell_{2},\ell_{3}$ that is not satisfied by the assignment. Then, none of these literals is represented by a task decomposed from a variable task. Further, the clause task $c$ can only be decomposed into at most 2 tasks with different actions from $\ell_{1}$, $\ell_{2}$, and $\ell_{3}$. Without loss of generality, let the decomposition of this clause task not produce a task with action $\ell_{1}$. This literal occurs in exactly one other clause and the corresponding clause task can produce at most one task with action $\ell_{1}$. This contradicts having 2 occurrences of action $\ell_{1}$ in the plan. Thus, all the clauses are satisfied.
Observe that the created task network has depth ${C_{d}}=1$ and each compound task can be decomposed into one of at most ${C_{c}}=7$ different task networks which each contain at most ${C_{s}}=2$ tasks. ∎
Theorem 18.
Plan Validation and Action Executability are NP-hard even if ${C_{c}}=7$, ${C_{d}}=2$, ${C_{\#}}=1$, $\prec^{+}=\emptyset$, and $\operatorname{prec}(a)$=$\operatorname{del}(a)$=$\operatorname{add}(a)$$\emptyset$ for every $a\in A$.
Proof.
Consider the task network $tn$ constructed in the proof of Theorem 17. Build a new initial task network $tn^{\prime}=(T={\{t\}},\prec^{+}=\emptyset,\alpha)$, and let $(\alpha(t),tn)\in M$. Then, ${C_{\#}}=1$, ${C_{d}}=2$, ${C_{c}}=7$, and after decomposing $t$ into $tn$, the reduction works exactly as in the proof of Theorem 17. ∎
Theorem 19.
Plan Validation and Action Executability are NP-hard even if ${C_{c}}=7$, ${C_{\#}}=1$, ${C_{s}}=3$, $\prec^{+}=\emptyset$, and $\operatorname{prec}(a)$=$\operatorname{del}(a)$=$\operatorname{add}(a)$$\emptyset$ for every $a\in A$.
Proof.
We adapt our construction used in the proof of Theorem 17. Let the 3-SAT-(2,2) instance have variables $x_{1},\ldots,x_{n}$ and clauses $C_{1},\ldots,C_{m}$. Once more, let the set of actions $A={\{x_{1},\overline{x_{1}},\ldots,x_{n},\overline{x_{n}}\}}$ contain an action for each literal, and let there be a compound task for each variable and clause, i.e., ${\cal C}={\{X_{1},\ldots,X_{n},C_{1},\ldots,C_{m}\}}$. For each $i\in[n-1]$, the variable task $X_{i}$ can be decomposed into one of two subnetworks. Both these subnetworks contain the compound task $X_{i+1}$, and one of them contains two tasks with action $x_{i}$, while the other contains two tasks with action $\overline{x_{i}}$. The decomposition for $X_{n}$ is similar, but instead of a compound task $X_{n+1}$, it adds the compound task $C_{1}$ in both subnetworks. Just like before, the compound tasks for clauses can be decomposed into one of the 7 subnetworks representing proper subsets of the literals in the clause. For $i\in[m-1]$, they additionally contain the compound task $C_{i+1}$. The initial network now contains only a single compound task $X_{1}$, giving ${C_{\#}}=1$. As before, each compound task can be decomposed into one of at most ${C_{c}}=7$ different task networks which, in this case, each contain at most ${C_{s}}=3$ tasks. However, the decomposition depth is no longer bounded by a constant. For Plan Validation, let the input plan $p$ be any sequence that contains each literal exactly twice. For Action Executability, let $S$ contain each literal exactly twice. The correctness of the reduction follows from the same arguments that were used in the proof of Theorem 17. ∎
We conclude this section with an analogous intractability result complementing Theorem 16, even if we consider the “classical” partial order width as our stable measure, as opposed to the generalized variant considered here:
Theorem 20.
Plan Existence, Action Executability, and State Reachability are NP-hard in task networks where $\prec^{+}$ describes a total order on $T$ and all subnetworks in $M$ are total orders even if ${C_{c}}=2$, ${C_{d}}=1$, and ${C_{s}}=1$. The same holds if ${C_{c}}=2$, ${C_{d}}=2$, and ${C_{\#}}=1$, or ${C_{c}}=2$, ${C_{\#}}=1$, and ${C_{s}}=2$.
Proof.
@@ 5 Hierarchical Task Networks
We reduce from $k$-Multicolored-Clique on a graph $G=(V,E)$, which is NP-hard for unbounded $k$ Fellows et al. (2009). Let $tn(p,d,a)$ with $p,d,a\subseteq F$ refer to a task network with a single task $t$ with action $\alpha(t)$ such that $\operatorname{prec}(\alpha(t))=p$, $\operatorname{del}(\alpha(t))=d$, and $\operatorname{add}(\alpha(t))=a$. For each vertex $v_{i}$, $i\in|V|$, create a compound task $V_{i}$, and, for each edge $e_{i}$, $i\in|E|$, create a compound task $E_{i}$. Create a primitive task $t_{g}$ with action $a_{g}$ such that $\operatorname{prec}(a_{g})={\{\text{EDGE}_{j,j^{\prime}}\mid j,j^{\prime}\in[k],j<j^{\prime}\}}$, $\operatorname{del}(a_{g})=\emptyset$, and $\operatorname{add}(a_{g})={\{\text{GOAL}\}}$. The initial network $tn$ contains exactly these tasks in a total order where $(V_{i},V_{i+1})\in\prec$ for all $i\in[|V|-1]$, $(V_{|V|},E_{1})\in\prec$, $(E_{i},E_{i+1})\in\prec$ for all $i\in[|E|-1]$, and $(E_{|E|},t_{g})\in\prec$. Each vertex and edge compound task can be decomposed into a primitive task with no effect, that is, $(V_{i},tn(\emptyset,\emptyset,\emptyset))\in M$ for all $i\in[|V|]$, and $(E_{i},tn(\emptyset,\emptyset,\emptyset))\in M$ for all $i\in[|E|]$. This corresponds to not selecting respective vertices and edges as part of the clique. Further, a vertex can be chosen for color $j$ if no other vertex has been chosen for this color (represented by having $\text{COLOR}_{j}$ in the state). Thus, $(V_{i},tn({\{\text{COLOR}_{j}\}},{\{\text{COLOR}_{j}\}},{\{v_{i}\}}))\in M$ for $i\in[|V|]$ and the color $j$ of the vertex $v_{i}$. An edge can be chosen as the $j^{\text{th}}$ edge in the clique if both its vertices are chosen. Thus, for all $i\in[|E|]$, if $e_{i}=(v,v^{\prime})$, $v$ has color $j$, and $v^{\prime}$ has color $j^{\prime}$, then $(E_{i},tn({\{v,v^{\prime}\}},\emptyset,{\{\text{EDGE}_{j,j^{\prime}}\}}))\in M$. The construction is concluded by setting the initial state $s_{0}={\{\text{COLOR}_{i}\mid i\in[k]\}}$, setting $S={\{a_{g}\}}$ for Action Executability, and setting $s_{g}={\{\text{GOAL}\}}$ for State Reachability. The construction clearly takes polynomial time.

### v2

@@ Appendix A Appendix: Proof Details
For an Action Executability instance, first test whether Reduction Rule R0 can be applied, and if yes we use it to terminate. Let $k$ be the number of states and $w$ the generalized partial order width of the input instance. Set $U$ to be the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ can be decomposed into $w$ chains $c_{1},\ldots,c_{w}$. This takes $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin [2003]. Let $A={\{a_{1},\ldots,a_{|A|}\}}$ and let $E={\{e_{1},\ldots,e_{|E|}\}}$ denote the set of equivalence classes among all actions, where action equivalence is defined as in the proof of Theorem 4. For each state $s$ in the state transition graph, let $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ be a boolean variable that is TRUE if and only if there is a sequence of tasks that respects $\prec^{+}$, is executable from $s_{0}$, yields state $s$ and employs exactly the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and contains exactly $r_{i}$ tasks from $U$ with action equivalence class $e_{i}$, for all $i\in[|E|]$.
@@
Assuming the variable is set correctly for all possible assignments, the instance can be decided as follows. For State Reachability, we have a yes-instance if and only if there are $h_{1},\ldots,h_{w},r_{1},\ldots,r_{w}$ such that $R_{s_{g}}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ is TRUE. For Action Executability, for a variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ let $c(a)$ denote the number of tasks in $V$ with action $a$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. We have a yes-instance if and only if there is a TRUE variable such that, for each action $a$ there are at least $\min(a)-c(a)$ tasks of action $a$ in $U$ and for each action equivalence class $e_{i}\in E$, we have $\sum_{a\in e_{i}}\max{\{0,\min(a)-c(a)\}}\leq r_{i}$.
We compute all variable values in a dynamic programming manner. Initialize all values to FALSE except for $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$, which is initialized to TRUE. Whenever a variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ that was FALSE before is set to TRUE (including the initial one), also perform the following update step. Let $T^{\prime}\subseteq T$ be such that $t\in T^{\prime}$ if, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Consider all the states $s^{\prime}$ that are adjacent to $s$ in the state transition graph ${\cal G}$, including $s$ itself if it has a self-loop. If there is $j\in[w]$ such that the $(h_{j}+1)^{\text{th}}$ task of the chain $c_{j}$ is in $T^{\prime}$, has action $a_{i}$, and $a_{i}$ connects $s$ and $s^{\prime}$ in ${\cal G}$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ to TRUE. This corresponds to taking the next task from a chain and adding it to a given sequence. Another update rule corresponds to adding a task from $U$ instead. If there is an action equivalence class $e_{i}$ that connects $s$ and $s^{\prime}$ in ${\cal G}$, then define the following. Let $T_{i}={\{t\in T\mid\alpha(t)\in e_{i}\}}$ and let $v_{i}$ denote the number of tasks in $V\cap T_{i}$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. If $r_{i}-v_{i}<|T_{i}\cap U|$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{i-1},r_{i}+1,r_{i+1},\ldots,r_{|E|}]$ to TRUE. Once no more variables are updated and set to TRUE, the computation is complete.
We show that the algorithm sets all variables correctly by an inductive argument over $z=\sum_{i=1}^{w}h_{i}+\sum_{i=1}^{|E|}r_{i}$. When one TRUE variable sets another variable to TRUE, this increases this sum by exactly 1. For $z=0$, the algorithm sets all variables correctly as the only way to build an empty plan is to take no tasks from any chain and remain in the initial state, and thus, the only TRUE variable in this case is $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$. Now, assume that the algorithm correctly assigned all variables up to some arbitrary, but fixed $z$. Consider a variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{w}]$ with $\sum_{i=1}^{w}h_{i}+\sum_{i=1}^{|E|}r_{i}=z+1$.
If the algorithm sets the variable to be TRUE, we distinguish between two cases. First, assume it does so starting from some TRUE variable $R_{s^{\prime}}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{i-1},r_{i}-1,r_{i+1},\ldots,r_{|E|}]$ with $i\in[|E|]$. Then, $a_{i}$ is executable in state $s^{\prime}$ and transforms it into $s$. Further, due to the induction hypothesis, there is an executable task sequence that uses the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and consists of $r_{\ell}$ tasks of action $a_{\ell}$, for each $\ell\in[|A|]\setminus{\{i\}}$, and $r_{i}-1$ tasks of action $a_{i}$. Thus, it contains $r_{i}-v_{i}$ tasks in $U$ with action equivalence class $e_{i}$. As there are more than $r_{i}-v_{i}$ tasks of that equivalence class in $U$ and their action is executable in the current state, one of the remaining tasks of action $a_{i}$ can be appended to the sequence. The resulting sequence is executable from $s_{0}$, respects $\prec^{+}$, and uses the same number of tasks of each chain and action, except one more task with action $a_{i}$. Thus, $R_{s^{\prime}}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{w}]$ is indeed TRUE.
In the other case, it sets the variable to TRUE starting from some $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},h_{w}][r_{1},\ldots,r_{i-1},r_{i}-1,r_{i+1},\ldots,r_{|E|}]$ being TRUE with $j\in[w]$, $i\in[|E|]$. Then, due to the induction hypothesis, there is a task sequence corresponding to an executable action sequence that uses the first $h_{\ell^{\prime}}$ tasks of each chain $c_{\ell^{\prime}}$, for each $\ell^{\prime}\in[w]\setminus\{j\}$, but only uses the first $h_{j}-1$ tasks of the chain $c_{j}$, and it consists of $r_{\ell}$ tasks of action equivalence class $r_{\ell}$, for each $\ell\in[|E|]\setminus{\{i\}}$, and $r_{i}-1$ tasks of action equivalence class $e_{i}$. Further, the $h_{j}^{\text{th}}$ task of chain $h_{j}$ has action equivalence class $e_{i}$, transforms state $s^{\prime}$ into $s$, and appending it to the sequence respects $\prec^{+}$. Thus, in this case it is also correct to set the variable to TRUE.
For the other direction, suppose the variable is supposed to be TRUE and let $\bar{t}$ be a task sequence witnessing its correctness. Consider the variable corresponding to the state, and values of the $h_{i}$ and $r_{j}$ after executing all but the last element of $\bar{t}$. As this shortened task sequence is a prefix of $\bar{t}$, its corresponding variable is supposed to be TRUE and the algorithm correctly identifies this by the induction hypothesis. When the algorithm sets this variable to be TRUE, in the following update steps it considers all possible next tasks in the sequence, including the last element of $\bar{t}$, and hence correctly sets the value of that variable to be TRUE.
@@
Computing the equivalence classes takes $\mathcal{O}\!\left(|{\cal I}|^{3}\right)$ time. As there are $|E|\leq(k+1)^{k}$ equivalence classes, at most $\mathcal{O}\!\left(|T|^{w}\cdot k\cdot|T|^{(k+1)^{k}}\right)$ variables are considered. For each of these, computing all $v_{i}$ takes $\mathcal{O}\!\left(|T|\right)$ time by iterating through all chains and it can update at most $w+|E|$ variables. Further, it takes $\mathcal{O}\!\left(|T|\right)$ time to check whether a variable yields a yes-instance. Hence, the algorithm takes ${\mathcal{O}\!\left(|T|^{w+(k+1)^{k}}\cdot k\cdot(|T||+w+|E|)+|V|^{2.5}+|{\prec}|\right)}={\mathcal{O}\!\left(|T|^{w+(k+1)^{k}+1}\cdot k\right)}$ time. ∎
Proof of Theorem 9.
@@
We reduce from the NP-hard Multicolored-Clique problem, which asks whether a given $k$-partite graph $G=(V,E)$ contains a $k$-clique Fellows et al. [2009]. Let $tn(p,d,a)$ with $p,d,a\subseteq F$ refer to a task network with a single task $t$ with action $\alpha(t)$ such that $\operatorname{prec}(\alpha(t))=p$, $\operatorname{del}(\alpha(t))=d$, and $\operatorname{add}(\alpha(t))=a$. For each vertex $v_{i}$, $i\in|V|$, create a compound task $V_{i}$, and, for each edge $e_{i}$, $i\in|E|$, create a compound task $E_{i}$. Create a primitive task $t_{g}$ with action $a_{g}$ such that $\operatorname{prec}(a_{g})={\{\text{EDGE}_{j,j^{\prime}}\mid j,j^{\prime}\in[k],j<j^{\prime}\}}$, $\operatorname{del}(a_{g})=\emptyset$, and $\operatorname{add}(a_{g})={\{\text{GOAL}\}}$. The initial network $tn$ contains exactly these tasks in a total order where $(V_{i},V_{i+1})\in\prec$ for all $i\in[|V|-1]$, $(V_{|V|},E_{1})\in\prec$, $(E_{i},E_{i+1})\in\prec$ for all $i\in[|E|-1]$, and $(E_{|E|},t_{g})\in\prec$. Each vertex and edge compound task can be decomposed into a primitive task with no effect, that is, $(V_{i},tn(\emptyset,\emptyset,\emptyset))\in M$ for all $i\in[|V|]$, and $(E_{i},tn(\emptyset,\emptyset,\emptyset))\in M$ for all $i\in[|E|]$. This corresponds to not selecting respective vertices and edges as part of the clique. Further, a vertex can be chosen for color $j$ if no other vertex has been chosen for this color (represented by having $\text{COLOR}_{j}$ in the state). Thus, $(V_{i},tn({\{\text{COLOR}_{j}\}},{\{\text{COLOR}_{j}\}},{\{v_{i}\}}))\in M$ for $i\in[|V|]$ and the color $j$ of the vertex $v_{i}$. An edge can be chosen as the $j^{\text{th}}$ edge in the clique if both its vertices are chosen. Thus, for all $i\in[|E|]$, if $e_{i}=(v,v^{\prime})$, $v$ has color $j$, and $v^{\prime}$ has color $j^{\prime}$, then $(E_{i},tn({\{v,v^{\prime}\}},\emptyset,{\{\text{EDGE}_{j,j^{\prime}}\}}))\in M$. The construction is concluded by setting the initial state $s_{0}={\{\text{COLOR}_{i}\mid i\in[k]\}}$, setting $S={\{a_{g}\}}$ for Action Executability, and setting $s_{g}={\{\text{GOAL}\}}$ for State Reachability. The construction takes polynomial time.